Hi I need answers for this problem.
3a+4b=2
-3a+b=23

Answer by lwsshak3(11628)

Hi I need answers for this problem.
3a+4b=2
-3a+b=23
add
5b=25
b=5
..
3a=-4b+2=-20+2=-18
3a=18
a=-6